ROTATIONAL ANALYSIS OF THE LASER INDUCED FLUORESCENCE EXCITATION SPECTRUM OF JET-COOLED CF_{3}O AND CF_{3}S

\underline{MIN-CHIEH YANG}, J. M. WILLIAMSON, AND TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

The high resolution, rotationally resolved laser induced fluorescence spectra for the Ã ^{2}A_{1} ↔ X̃ ^{2}E transition of CF_{3}O and CF_{3}S were recorded. The Ã electronic state symmetric vibrational bands 0_{0}^{0} and 3_{0}^{1} and the asymmetric bands 5_{0}^{1} and 6_{0}^{1} were analyzed for CF_{3}O, while for CF_{3}S the symmetric bands 0_{0}^{0}, 1_{0}^{1}, 2_{0}^{1} and 3_{0}^{1}, as well as the asymmetric bands 6_{0}^{1}, 3_{0}^{1}5_{0}^{1}, and 3_{0}^{1}6_{0}^{1} were analyzed. The results of the analysis show that the ground states of both CF_{3}O and CF_{3}S have C_{3v} symmetry which is consistent with a dynamic Jahn-Teller effect in these radicals. In the asymmetric 5_{0}^{1} band of CF_{3}O, both perpendicular and parallel type transitions were observed while for the other asymmetric bands, only perpendicular type transitions were observed. The molecular parameters and the details of the analysis of all these bands will be presented.